# Dynamical tide modified Roche limit in eccentric, asynchronous binaries

### v1

@@ Abstract
The Roche limit, or the threshold separation within which a celestial object (the donor) $M$ cannot remain in a stable configuration due to a companion’s tidal field, has been well established when $M$ is in hydrostatic equilibrium and has synchronous rotation in a circular orbit. However, limited analyses exist considering corrections to the Roche limit due to hydrodynamical effects. We fill in the gap by providing a general theoretical framework involving nonlinear hydrodynamics. We consider both exact nonlinear equations derived from an affine model describing incompressible ellipsoids and series-expanded ones that can be calculated for realistic stars and planets. Our formulation addresses the Roche problem in generic orbits and synchronization levels of $M$, and fully accounts for the history-dependent hydrodynamical effects. We show that as the orbital eccentricity increases, fluid instability is more likely to develop at the pericenter due to the increased dynamical tide that accumulates over multiple orbits. When $M$ moves in a highly eccentric orbit (with eccentricity around 0.9) and the damping of the fluid is small, the threshold pericenter separation at which mass loss from $M$ can occur can be at least 30% higher than the value predicted for a circular orbit with hydrostatic equilibrium. If only a single passage is considered, however, the threshold separation is 20% smaller than the static limit. The nonlinear interaction at each pericenter passage can also trigger a chaotic fluid evolution inside $M$ even with moderate eccentricities, complementing previous studies of chaotic tides caused by random propagation phases. Our work has broad implications for interacting binaries in eccentric orbits, including migrating gaseous exoplanets, repeated partial tidal disruption events, and more.
@@
software: NumPy (C. R. Harris et al., 2020), SciPy (P. Virtanen et al., 2020), Matplotlib (J. D. Hunter, 2007), Numba (S. K. Lam et al., 2015), Mathematica (Wolfram Research, Inc., 2025).
1 Introduction
@@
The Roche model has long been the foundational framework for studying mass loss and transfer in interacting binary systems. Of particular importance is the Roche limit, a critical separation within which a donor object (with mass $M$ and radius $R$) can no longer hold itself but will lose mass due to the tidal force of the companion (treated as a point particle with mass $M^{\prime}$ here). Many applications of the Roche model rely on the fit by P. P. Eggleton (1983), which assumes that the donor is in hydrostatic equilibrium–a valid approximation for evolved circular binaries but one that may break down in eccentric systems, particularly when eccentricity is high and the donor has not reached synchronization.
@@
Highly eccentric binaries can form in various environments. One such example lies in the realm of exoplanets. The formation of hot Jupiters (HJs) and hot Neptunes (HNs) at separations as close as $0.05\,{\rm AU}$ to their host stars is one of the longest outstanding puzzles in exoplanetary science (M. Mayor & D. Queloz, 1995; R. I. Dawson & J. A. Johnson, 2018). High-eccentricity migration of HJ and HN is a promising explanation (D. Fabrycky & S. Tremaine, 2007): A giant planet initially forms with a semimajor axis of $\mathcal{O}({\rm AU})$ and is dynamically excited onto a highly eccentric orbit. It then migrates inward due to tidal dissipation while the orbit circularizes. This picture naturally reproduces some key characteristics of HJs (R. I. Dawson & J. A. Johnson, 2018), including their highly inclined, sometimes retrograde orbits. While qualitatively successful, this channel still has large uncertainties on a quantitative level. For example, the pericenter separation required for inward migration can be dangerously close to tidal disruption of the planet, and a small variation in the disruption threshold can have a significant impact on the formation efficiency of HJs and HNs (H. Yu & F. Dai, 2024).
@@
Another example is found in the nucleus of a galaxy that hosts a massive black hole (MBH). In such regions, a stellar-mass binary may be drawn close to the MBH and disrupted by its tidal forces through the Hills mechanism, which ejects one star while leaving the other bound to the MBH on an eccentric orbit (J. G. Hills, 1988). If the bound star approaches the MBH closely enough, it may itself be disrupted, resulting in a tidal disruption event (TDE; M. J. Rees 1988; S. Komossa 2015). At slightly greater separations, the star may survive a full disruption by the MBH and repeat mass transfer over multiple orbits, causing repeated partial TDEs (rpTDEs; K. Hayasaki et al. 2013; E. R. Coughlin & C. J. Nixon 2019; M. Cufari et al. 2022; Z. Liu et al. 2024). Recently, a new class of recurring X-ray bursts, known as quasi-periodic eruptions (QPEs), has been observed and is associated with low-mass MBHs (G. Miniutti et al., 2019). The mechanism driving QPEs remains uncertain, yet similar to the production of TDEs (K. Maguire et al., 2020), several proposed models involve a star or white dwarf (WD) in an eccentric orbit around an MBH, undergoing periodic mass transfer episodes that power the QPEs (A. King, 2020, 2022; M. Wang et al., 2022; I. Linial & R. Sari, 2023). The recent work of S. Y. Lau & H. Yu (2025) demonstrates that both can be the result of a tidally-induced mass loss at different pericenter distances. For orbits with moderately high eccentricities ($0.90<e<0.99$) where the orbital period is short compared to the tidal dissipation timescale, the tide inside the WD builds up chaotically over multiple orbits. The tidal energy eventually approaches the stellar binding energy and is released via a mass ejection as the extended tidal bulge breaks. Neither the WD nor the orbit will be disrupted by a single break, allowing the process to repeat several times. This may lead to rpTDEs with time intervals of a few years, consistent with the observations (e.g., S. Campana et al. 2015; T. Wevers et al. 2023; G. Miniutti et al. 2023; J. J. Somalwar et al. 2025). Meanwhile, the model also shows the possibility of QPEs being sourced by tidal stripping at closer pericenter separations, where the tidal field is strong enough such that a single passage is sufficient to induce mass transfer, though this mechanism is challenged by certain observed features in some QPEs, like the alternating long-short recurrence time and brightness of bursts (G. Miniutti et al., 2019; M. Giustini et al., 2020; R. Arcodia et al., 2022; C. Zhou et al., 2024). The proposed mechanism by S. Y. Lau & H. Yu (2025) accounts for the dynamical nature of the tide over multiple orbits, improving previous studies of A. King (2022); M. Wang et al. (2022) based on the static Roche model that does not apply in eccentric orbits. Nonetheless, the threshold distance for the tidally induced mass transfer, either via multiple orbit buildup or a single passage, depends on a phenomenological parametrization that is not derived from fundamental hydrodynamics.
@@
Despite the importance of mass loss in eccentric binaries, illustrated by the examples above, theoretical advancements beyond the classic Roche model are still limited. Some pioneering work by J. F. Sepinsky et al. (2007) extended the Roche model to non-circular orbits with non-synchronized donors, identifying significant corrections to the classic fit of P. P. Eggleton (1983); however, their work was still based on a quasi-static approximation. In reality, the donor also exhibits a dynamical response, corresponding to oscillations of the eigenmodes (mainly the fundamental modes, or f-modes) excited at each pericenter passage (W. H. Press & S. A. Teukolsky, 1977). This component is known as the dynamical tide. The analysis of P. Diener et al. (1995), based on an affine model developed by B. Carter & J. P. Luminet (1983), accounted for the dynamical tide but focused on the hydrodynamical stability in a single passage (though they considered the effects due to varying the ellipsoid’s orientation). Unlike the equilibrium component, which depends solely on the instantaneous orbital configuration, the dynamical tide can accumulate over multiple orbits, allowing its amplitude to grow to a scale comparable to the unperturbed radius (R. A. Mardling, 1995; P. B. Ivanov & J. C. B. Papaloizou, 2004; Y. Wu, 2018; M. Vick & D. Lai, 2018; H. Yu et al., 2021). Such an accumulation is expected to play a critical role in the dynamics of mass transfer.
@@
The numerical study by J. Guillochon et al. (2011) examining the threshold separation at which a proto-HJ is disrupted by its host star strongly motivates the need to include DT in modeling mass transfer. According to the fit of P. P. Eggleton (1983), Roche lobe overflow should only occur when the planet’s pericenter separation $D_{p}$ reaches $D_{p}{\lesssim}2.0R_{t}$ with $R_{t}{=}(M^{\prime}/M)^{1/3}R$ the tidal radius. However, J. Guillochon et al. (2011) found that with high orbital eccentricity (e=0.9e{=}0.9), a planet is disrupted at a much greater pericenter separation of $D_{p}\simeq 2.7R_{t}$, $35\%$ greater than the separation from Eggleton’s fit. The discrepancy is further amplified when the results of J. F. Sepinsky et al. (2007) are used, which predict a Roche lobe $\sim 10\%$ greater than Eggleton’s fit (mainly due to the lack of centrifugal force), requiring an even smaller separation for the overflow to occur. Similarly, the critical single-passage separation of $D_{p}\simeq 1.7R_{t}$ from P. Diener et al. (1995) also cannot explain the numerical results of J. Guillochon et al. (2011).
@@
In this work, we will show that the dynamical tide is the needed ingredient to reconcile the discrepancy. For this, we will consider the nonlinear hydrodynamical equations following both the affine model of P. Diener et al. (1995) (see also B. Carter & J. P. Luminet 1983, 1985) and those derived by T. Van Hoolst (1994) following an order-by-order expansion (see also N. N. Weinberg et al. 2012; H. Yu et al. 2023). Our model is fully dynamical and makes no (quasi-)static assumptions. Consequently, it applies to arbitrary eccentricities of the orbit and asynchronicities of the donor (though we still restrict our analysis to aligned spins). On the other hand, when the perturbed fluid is in hydrostatic equilibrium, our model also reproduces the classic results obtained by S. Chandrasekhar (1963) (see also S. Chandrasekhar 1987 for a comprehensive summary) on homogeneous ellipsoids.
@@
The rest of the paper is organized as follows. We will first describe a toy model in §1.1 that captures the essence of our model and illustrates the significance of the dynamical tides. The nonlinear hydrodynamical model is introduced in §2, after which we will demonstrate its consistency with the analysis of S. Chandrasekhar (1963) in §2.1 under the static limit. A theoretical investigation on instantaneous fluid stability at the pericenter follows in §2.2. Then, in §2.3, a numerical experiment is conducted to address the build-up of the instability throughout multiple orbits. How the model applies to different astrophysical settings is described in §3. Lastly, we conclude in §4.
@@
1.1 A toy model
We use $\xi\in\mathbb{R}$ to denote a real fluid displacement in $M$ and $\epsilon>0$ a tidal perturbation. We will show later that the problem can be represented by a simple toy model
$\ddot{\xi}=\omega_{f}^{2}(\xi-I\epsilon-\kappa\xi^{2}),$ (1)
where $\omega_{f}^{2}$, $I$, and $\kappa$ are all positive constants. Physically, they correspond to the frequency squared of an eigenmode and the mode’s linear and nonlinear coupling coefficients. For the linear problem (ignoring the $\kappa\xi^{2}$ term), a static solution ($\ddot{\xi}=0$) can always be arranged by setting $\xi=I\epsilon$. With the nonlinear $\kappa\xi^{2}$ term, the right-hand side (RHS) of Eq. (1) must take the same sign as $\xi$ for the system to be oscillatory. We immediately see that with $\kappa\xi^{2}\leq 0$, the nonlinear term can lead to instability when $\xi>0$, and an increasing $\xi$ will make the system more vulnerable to instability. If we write $\xi\sim e^{\sigma t}$, at the instability threshold ($\sigma=0$), we need
$\kappa\xi^{2}+\xi-I\epsilon=0,$ (2)
@@ 1.1 A toy model
We thus obtain that $1/(4\kappa I)$ is the maximum perturbation allowed. This is the essence of the static Roche limit found in S. Chandrasekhar (1963). See later in §2.1.
To see the effect of the dynamical effects, we first consider $\xi\sim\cos(\varpi t)e^{\sigma t}$ with $\varpi^{2}\lesssim\omega_{f}^{2}$. For the threshold $\sigma=0$, we need
$\omega_{f}^{2}\kappa\xi^{2}+(\omega_{f}^{2}-\varpi^{2})\xi-\omega_{f}^{2}I\epsilon=0,$ (4)
@@
Compared to the static case, Eq. (3), including the dynamical effect makes instability more likely to occur as the maximum $\epsilon$ is smaller by a factor of $(\omega_{f}^{2}-\varpi^{2})^{2}/\omega_{f}^{4}$. This model will be relevant for the discussions in §2.2 about finite-frequency corrections to the equilibrium tide.
@@
The $\xi_{\rm dyn}$ term reduces (increases) the maximum $\epsilon$ a stable configuration can allow if $\xi_{\rm dyn}>0$ ($\xi_{\rm dyn}<0$). If a positive $\xi_{\rm dyn}$ has triggered some exponential growth in $\eta$, suddenly changing its sign will only make $\eta$ oscillatory but not damp away the growth. Therefore, for a sufficiently large $\xi_{\rm dyn}$ with a randomly fluctuating sign, the net effect will again make the instability more likely to occur. This model is relevant to the dynamical tide in §2.3.
@@
2 Nonlinear hydrodynamics
We focus mainly on a homogeneous, incompressible fluid object with a polytropic index $n=0$ and an adiabatic index $\Gamma\to\infty$ for the theoretical discussion, though our numerical investigations later in §2.3 will be extended to touch upon $n=1$ polytropes. This choice is for simplicity and to assist a direct comparison with the classical results obtained by S. Chandrasekhar (1963, 1987). Denote the Lagrangian displacement of a mass element $dM$ as $\boldsymbol{\xi}=\boldsymbol{r}-\boldsymbol{x}$ with $\boldsymbol{r}$ and $\boldsymbol{x}$ the perturbed and original position vectors, respectively. We further decompose $\boldsymbol{\xi}$ into eigenmodes in the configuration space, as

### v2

@@ Abstract
The Roche limit, or the threshold separation within which a celestial object (the donor) $M$ cannot remain in a stable configuration due to a companion’s tidal field, has been well established when $M$ is in hydrostatic equilibrium and has synchronous rotation in a circular orbit. However, limited analyses exist considering corrections to the Roche limit due to hydrodynamical effects. We fill in the gap by providing a general theoretical framework involving nonlinear hydrodynamics. We consider both exact nonlinear equations derived from an affine model describing incompressible ellipsoids and series-expanded ones that can be calculated for realistic stars and planets. Our formulation addresses the Roche problem in generic orbits and synchronization levels of $M$, and fully accounts for the history-dependent hydrodynamical effects. We show that as the orbital eccentricity increases, fluid instability is more likely to develop at the pericenter due to the increased dynamical tide that accumulates over multiple orbits. When $M$ moves in a highly eccentric orbit (with eccentricity around 0.9) and the damping of the fluid is small, the threshold pericenter separation at which mass loss from $M$ can occur can be at least 30% higher than the value predicted for a circular orbit with hydrostatic equilibrium. If only a single passage is considered, however, the threshold separation is 20% smaller than the static limit. The nonlinear interaction at each pericenter passage can also trigger a diffusive fluid evolution inside $M$ even with moderate eccentricities, complementing previous studies of diffusive tides caused by random propagation phases. Our work has broad implications for interacting binaries in eccentric orbits, including migrating gaseous exoplanets, repeated partial tidal disruption events, and more.
@@
software: NumPy (Harris et al., 2020), SciPy (Virtanen et al., 2020), Matplotlib (Hunter, 2007), Numba (Lam et al., 2015), Mathematica (Wolfram Research, Inc., 2025).
I Introduction
@@
The Roche model has long been the foundational framework for studying mass loss and transfer in interacting binary systems. Of particular importance is the Roche limit, a critical separation within which a donor object (with mass $M$ and radius $R$) can no longer hold itself but will lose mass due to the tidal force of the companion (treated as a point particle with mass $M^{\prime}$ here). Many applications of the Roche model rely on the fit by Eggleton (1983), which assumes that the donor is in hydrostatic equilibrium–a valid approximation for evolved circular binaries but one that may break down in eccentric systems, particularly when eccentricity is high and the donor has not reached synchronization.
@@
Highly eccentric binaries can form in various environments. One such example lies in the realm of exoplanets. The formation of hot Jupiters (HJs) and hot Neptunes (HNs) at separations as close as $0.05\,{\rm AU}$ to their host stars is one of the longest outstanding puzzles in exoplanetary science (Mayor and Queloz, 1995; Dawson and Johnson, 2018). High-eccentricity migration of HJ and HN is a promising explanation (Fabrycky and Tremaine, 2007): A giant planet initially forms with a semimajor axis of $\mathcal{O}({\rm AU})$ and is dynamically excited onto a highly eccentric orbit. It then migrates inward due to tidal dissipation while the orbit circularizes. This picture naturally reproduces some key characteristics of HJs (Dawson and Johnson, 2018), including their highly inclined, sometimes retrograde orbits. While qualitatively successful, this channel still has large uncertainties on a quantitative level. For example, the pericenter separation required for inward migration can be dangerously close to tidal disruption of the planet, and a small variation in the disruption threshold can have a significant impact on the formation efficiency of HJs and HNs (Yu and Dai, 2024).
@@
Another example is found in the nucleus of a galaxy that hosts a massive black hole (MBH). In such regions, a stellar-mass binary may be drawn close to the MBH and disrupted by its tidal forces through the Hills mechanism, which ejects one star while leaving the other bound to the MBH on an eccentric orbit (Hills, 1988). If the bound star approaches the MBH closely enough, it may itself be disrupted, resulting in a tidal disruption event (TDE; Rees 1988; Komossa 2015). At slightly greater separations, the star may survive a full disruption by the MBH and repeat mass transfer over multiple orbits, causing repeated partial TDEs (rpTDEs; Hayasaki et al. 2013; Coughlin and Nixon 2019; Cufari et al. 2022; Liu et al. 2024; Bandopadhyay et al. 2024, 2025; Yao and Quataert 2025). Recently, a new class of recurring X-ray bursts, known as quasi-periodic eruptions (QPEs), has been observed and is associated with low-mass MBHs (Miniutti et al., 2019). The mechanism driving QPEs remains uncertain, yet similar to the production of TDEs (Maguire et al., 2020), several proposed models involve a star or white dwarf (WD) in an eccentric orbit around an MBH, undergoing periodic mass transfer episodes that power the QPEs (King, 2020, 2022; Wang et al., 2022; Linial and Sari, 2023). The recent work of Lau and Yu (2025) demonstrates that both can be the result of a tidally-induced mass loss at different pericenter distances. For orbits with moderately high eccentricities ($0.90<e<0.99$) where the orbital period is short compared to the tidal dissipation timescale, the tide inside the WD builds up diffusively (i.e., as a random-walk process) over multiple orbits. The tidal energy eventually approaches the stellar binding energy and is released via a mass ejection as the extended tidal bulge breaks. Neither the WD nor the orbit will be disrupted by a single break, allowing the process to repeat several times. This may lead to rpTDEs with time intervals of a few years, consistent with the observations (e.g., Campana et al. 2015; Wevers et al. 2023; Miniutti et al. 2023; Somalwar et al. 2025). Meanwhile, the model also shows the possibility of QPEs being sourced by tidal stripping at closer pericenter separations, where the tidal field is strong enough such that a single passage is sufficient to induce mass transfer, though this mechanism is challenged by certain observed features in some QPEs, like the alternating long-short recurrence time and brightness of bursts (Miniutti et al., 2019; Giustini et al., 2020; Arcodia et al., 2022; Zhou et al., 2024). The proposed mechanism by Lau and Yu (2025) accounts for the dynamical nature of the tide over multiple orbits, improving previous studies of King (2022); Wang et al. (2022) based on the static Roche model that does not apply in eccentric orbits. Nonetheless, the threshold distance for the tidally induced mass transfer, either via multiple orbit buildup or a single passage, depends on a phenomenological parametrization that is not derived from fundamental hydrodynamics.
@@
Despite the importance of mass loss in eccentric binaries, illustrated by the examples above, theoretical advancements beyond the classic Roche model are still limited. Some pioneering work by Sepinsky et al. (2007) extended the Roche model to non-circular orbits with non-synchronized donors, identifying significant corrections to the classic fit of Eggleton (1983); however, their work was still based on a quasi-static approximation. In reality, the donor also exhibits a dynamical response, corresponding to oscillations of the eigenmodes (mainly the fundamental modes, or f-modes) excited at each pericenter passage (Press and Teukolsky, 1977). This component is known as the dynamical tide. The analysis of Diener et al. (1995), based on an affine model developed by Carter and Luminet (1983), accounted for the dynamical tide but focused on the hydrodynamical stability in a single passage (though they considered the effects due to varying the ellipsoid’s orientation). Unlike the equilibrium component, which depends solely on the instantaneous orbital configuration, the dynamical tide can accumulate over multiple orbits, allowing its amplitude to grow to a scale comparable to the unperturbed radius (Mardling, 1995; Ivanov and Papaloizou, 2004; Wu, 2018; Vick and Lai, 2018; Yu et al., 2021). Such an accumulation is expected to play a critical role in the dynamics of mass transfer.
@@
The numerical study by Guillochon et al. (2011) examining the threshold separation at which a proto-HJ is disrupted by its host star strongly motivates the need to include DT in modeling mass transfer. According to the fit of Eggleton (1983), Roche lobe overflow should only occur when the planet’s pericenter separation $D_{p}$ reaches $D_{p}{\lesssim}2.0R_{t}$ with $R_{t}{=}(M^{\prime}/M)^{1/3}R$ the tidal radius. However, Guillochon et al. (2011) found that with high orbital eccentricity (e=0.9e{=}0.9), a planet is disrupted at a much greater pericenter separation of $D_{p}\simeq 2.7R_{t}$, $35\%$ greater than the separation from Eggleton’s fit. The discrepancy is further amplified when the results of Sepinsky et al. (2007) are used, which predict a Roche lobe $\sim 10\%$ greater than Eggleton’s fit (mainly due to the lack of centrifugal force), requiring an even smaller separation for the overflow to occur. Similarly, the critical single-passage separation of $D_{p}\simeq 1.7R_{t}$ from Diener et al. (1995) also cannot explain the numerical results of Guillochon et al. (2011).
@@
In this work, we will show that the dynamical tide is the needed ingredient to reconcile the discrepancy. For this, we will consider the nonlinear hydrodynamical equations following both the affine model of Diener et al. (1995) (see also Carter and Luminet 1983, 1985) and those derived by Van Hoolst (1994) following an order-by-order expansion (see also Weinberg et al. 2012; Yu et al. 2023). Our model is fully dynamical and makes no (quasi-)static assumptions. Consequently, it applies to arbitrary eccentricities of the orbit and asynchronicities of the donor (though we still restrict our analysis to aligned spins). On the other hand, when the perturbed fluid is in hydrostatic equilibrium, our model also reproduces the classic results obtained by Chandrasekhar (1963) (see also Chandrasekhar 1987 for a comprehensive summary) on homogeneous ellipsoids.
@@
The rest of the paper is organized as follows. We will first describe a toy model in §I.1 that captures the essence of our model and illustrates the significance of the dynamical tides. The nonlinear hydrodynamical model is introduced in §II, after which we will demonstrate its consistency with the analysis of Chandrasekhar (1963) in §II.1 under the static limit. A theoretical investigation on instantaneous fluid stability at the pericenter follows in §II.2. Then, in §II.3, a numerical experiment is conducted to address the build-up of the instability throughout multiple orbits. How the model applies to different astrophysical settings is described in §III. Lastly, we conclude in §IV.
@@
I.1 A toy model
The appearance of the Roche limit can be schematically shown with a toy model:
$\ddot{\xi}=\omega_{f}^{2}(\underbrace{-\xi}_{\text{linear oscillator}}+\underbrace{I\epsilon}_{\text{linear tide}}+\underbrace{\kappa\xi^{2}}_{\text{combined nonlinear corrections}}),$ (1)
where $\xi\in\mathbb{R}$ denotes the fluid displacement due to a large-scale perturbation (i.e., the f-mode of the donor whose eigenfrequency is denoted by $\omega_{f}$). The first term on the right hand side (RHS) describes the linear restoration of the fluid to the equilibrium, the $I\epsilon$ term the linear tidal drive, with $I>0$ a coupling constant, $\epsilon=M^{\prime}/D^{3}$ the instantaneous tidal field exerted by a companion (the accretor) with mass $M^{\prime}$, and $D$ the instantaneous binary separation. Lastly, the $\kappa\xi^{2}$ term represents a combination of nonlinear effects in both the internal fluid response and the external tidal drive. Theoretical calculation predicts $\kappa>0$ when $I>0$ (see Appendix A). For the linear problem (ignoring the $\kappa\xi^{2}$ term), a static solution ($\ddot{\xi}=0$) can always be arranged by setting $\xi=I\epsilon$. With the nonlinear $\kappa\xi^{2}$ term, the RHS of Eq. (1) must take the opposite (same) sign as $\xi$ for the system to be oscillatory (grow exponentially). We immediately see that with $\kappa\xi^{2}\geq 0$, the nonlinear term can lead to instability when $\xi>0$, and an increasing $\xi$ will make the system more vulnerable to instability. This is consistent with the standard Roche lobe overflow picture where mass loss occurs mainly at the L1 and L2 Lagrangian points, corresponding to locations of maximum stretch of the donor. If we assume the hydrostatic limit and write $\xi\sim e^{\sigma t}$, at the instability threshold ($\sigma=0$), we need
$\kappa\xi^{2}-\xi+I\epsilon=0,$ (2)
@@ I.1 A toy model
We thus obtain that $1/(4\kappa I)$ is the maximum perturbation allowed. This is the essence of the static Roche limit found in Chandrasekhar (1963). See later in §II.1. Alternatively, Eq. (1) can also be viewed as a particle with displacement $\xi$ in an effective potential $U_{\rm eff}(\xi)$, with $\ddot{\xi}=-dU_{\rm eff}/d\xi$ and
$U_{\rm eff}(\xi)=\underbrace{\frac{1}{2}\omega_{f}^{2}\xi^{2}-\frac{1}{3}\kappa\omega^{2}_{f}\xi^{3}}_{\text{internal energy + self-gravity}}\underbrace{-I\epsilon\xi.}_{\text{external interaction}}$ (4)
As the donor is sufficiently deformed, the total energy of the internal fluid can decrease (see Appendix A), creating a local maximum in the effective potential. Going beyond the local maximum will then lead to a runaway in $\xi$ (Wu, 1998). [^1] In numerical experiments to be presented in §II.3, we confirm the presence of the instability by solving the fully nonlinear equations derived in Appendix A. While previous studies (e.g., Wu 1998) suggested that the four-wave interaction (the $\propto\xi^{4}$ term) in the effective potential could prevent the instability, the stabilization effect is eventually taken over by even higher-order corrections. When considering the f-modes, the full stability threshold is similar to that obtained at the three-wave order, as presented in the toy model here.
To see the effect of the dynamical effects, we first consider $\xi\sim\cos(\varpi t)e^{\sigma t}$ with $\varpi^{2}\lesssim\omega_{f}^{2}$. Plugging it back to Eq. (1) and consider the threshold case, $\sigma=0$, we have
$\omega_{f}^{2}\kappa\xi^{2}-(\omega_{f}^{2}-\varpi^{2})\xi+\omega_{f}^{2}I\epsilon=0,$ (5)
@@
Compared to the static case of Eq. (3), including the dynamical effect makes instability more likely to occur as the maximum $\epsilon$ is smaller by a factor of $(\omega_{f}^{2}-\varpi^{2})^{2}/\omega_{f}^{4}$. This model will be relevant for the discussions in §II.2 about finite-frequency corrections to the equilibrium tide, relevant mainly to asynchronous donors in mildly eccentric binaries.
@@
The $\xi_{\rm dyn}$ term reduces (increases) the maximum $\epsilon$ a stable configuration can allow if $\xi_{\rm dyn}>0$ ($\xi_{\rm dyn}<0$). If a positive $\xi_{\rm dyn}$ has triggered some exponential growth in $\eta$, suddenly changing its sign will only make $\eta$ oscillatory but not damp away the growth. Therefore, for a sufficiently large $\xi_{\rm dyn}$ with a randomly fluctuating sign, the net effect will again make the instability more likely to occur. This model is relevant to the dynamical tide in a donor moving in a highly eccentric orbit in §II.3.
@@
II Nonlinear hydrodynamics
We focus mainly on a homogeneous, incompressible fluid object with a polytropic index $n=0$ and an adiabatic index $\Gamma\to\infty$ for the theoretical discussion, though our numerical investigations later in §II.3 will be extended to touch upon $n=1$ polytropes. This choice is for simplicity and to assist a direct comparison with the classical results obtained by Chandrasekhar (1963, 1987). Denote the Lagrangian displacement of a mass element $dM$ as $\boldsymbol{\xi}=\boldsymbol{r}-\boldsymbol{x}$ with $\boldsymbol{r}$ and $\boldsymbol{x}$ the perturbed and original position vectors, respectively. We further decompose $\boldsymbol{\xi}$ into eigenmodes in the configuration space, as